# explExample

### v2

@@ 5.2 A time-layered approximation of the rarefied characteristics
(5.7) ,:${\vdash},\,{\dashv}:\{(x,t)\mid x\in\Upsilon_{t},t\in\Omega_{t}\}\to\widehat{\Omega}_{x},$
@@
Since $\widehat{X}(\cdot,t)$ is continuous and it is constant in the interval $[{\vdash}(x,t),{\dashv}(x,t)]$,
@@
(5.10) $\left\lvert{\vdash}(\widehat{\Omega}_{x},t)\right\rvert=\left\lvert{\dashv}(\widehat{\Omega}_{x},t)\right\rvert\leq\left\lvert\{x\in\widehat{\Omega}_{x}:u_{0}(x_{-})\neq u_{0}(x_{+})\}\right\rvert,$
where the RHS is the number of jumps in the initial condition $u_{0}$. For such an initial condition, the number of jumps does not increase [76].
@@
[…]
@@
[…]
@@
(5.13) $v_{k}(x):=\left\{\begin{aligned} (F^{\prime}\circ\hat{u}_{0})(x)&&\text{ if }\lambda(x)>t,\\ (F^{\prime}\circ\hat{u}_{0})({\vdash}(x,\,t_{k}))&&\text{ if }\lambda(x)\leq t.\end{aligned}\right.$
[…]
@@
Note that the upper bound (5.14) with the TV norm $\left\lvert{u_{0}}\right\rvert_{TV(\Omega_{x})}$ obtained from the last term in (B.25) is not sharp, and can be replaced by $2\lVert{u_{0}}\rVert_{L^{\infty}(\Omega_{x})}$ because
(5.16) $\left\lvert\overline{u}_{\text{loc}}-\underline{u}_{\text{loc}}\right\rvert\leq 2\lVert{u_{0}}\rVert_{L^{\infty}([{\vdash}(x_{0},t),{\dashv}(x_{0},t)])},$
but we will keep the TV norm here, as doing so this will not affect the final results of this work.
Suppose we let $(t_{k})_{k=1}^{K}$ be a uniform grid over time domain $\Omega_{t}$,
@@
See Figure 8 for an illustration of this approximation.
@@ 5.3 The spatial discretization of time-layered approximation
In the following, we will make use of the estimates (LABEL:eq:hXZk_error,eq:ZkTV) in Lemma 5.1, and the uniform time grid $t_{k}$ with grid size $K$ (5.17).
Because of the TV bound (5.15), there is an adaptive grid $(z_{i})_{i=1}^{K}$ over $\widehat{\Omega}_{x}$ on which we have an approximation $\barwedge_{k}(\cdot,t)\in C(\widehat{\Omega}_{x}\times\Omega_{t})\cap P_{1}(\widehat{\Omega}_{x}\times\Omega_{t})$ of $\Xi_{k}(\cdot,t)$ satisfying
(5.20) $\lVert{\Xi_{k}(\cdot,t)-\barwedge_{k}(\cdot,t)}\rVert_{L^{\infty}(\widehat{\Omega}_{x})}\leq\frac{1}{K}\left(1+C_{u_{0},T,F}\right).$
@@
(5.21) $\barwedge_{k}(z_{i})=\left\{\begin{aligned} \widehat{X}_{0}(z_{i},t_{k})&\quad\text{ for }z_{i}\in\widehat{\Omega}_{x}\setminus\Upsilon_{t_{k}},\\ \widehat{X}(z_{i},t_{k})&\quad\text{ for }z_{i}\in\Upsilon_{t_{k}}.\end{aligned}\right.$
@@
(5.22) $\eta_{k}(z_{i})=\frac{1}{t_{k+1}}(\barwedge_{k+1}(z_{i})-\hat{I}(z_{i}))-\frac{1}{t_{k}}(\barwedge_{k}(z_{i})-\hat{I}(z_{i}))\quad i=0,1,...\,,K-1.$
Then $\eta_{k}$ is nonzero only at the grid points $(z_{i})_{i=1}^{K}$ that lie in $\Upsilon_{t_{k+1}}\cup\Upsilon_{t_{k}}$. See Figure 9 for an illustration.
@@
Next, let us set the function $\barwedge:\widehat{\Omega}_{x}\times\Omega_{t}\to\mathbb{R}$ which is continuous piecewise linear in space but only piecewise linear in time. Let $\barwedge_{K}(x,t):=\hat{I}(x)+\frac{t}{T}(\barwedge_{K}(x)-\hat{I}(x))$
@@
$\leq\lVert{\widehat{X}(\cdot,t)-\Xi_{k}(\cdot,t)}\rVert_{L^{\infty}(\widehat{\Omega}_{x})}+\lVert{\Xi_{k}(\cdot,t)-\barwedge_{k}(\cdot)}\rVert_{L^{\infty}(\widehat{\Omega}_{x})}+\lVert{\barwedge_{k}(\cdot)-\barwedge(\cdot,t)}\rVert_{L^{\infty}(\widehat{\Omega}_{x})}$
@@
[…]
@@ Appendix A Glossary of definitions and relations
Formulations of the entropy solution (Section 2).
@@
    $\widehat{X}_{0}$ rarefied chars. with multivalued inverse
    (2.26) $\widehat{X}_{0}(x,t):=\hat{I}(x)+t(F^{\prime}\circ\hat{u}_{0})(x)$
@@
    $\rho$ threshold function (2.31)
@@
Low rank neural representation (LRNR) (Section 3).
@@
Transported subspaces, inverse-bias trick, LRNR approximation of classical solutions (Section 4).
@@
    LRNR ($L=2,r=3$) approximation $h(t)\in H_{r}$ to the classical solution (Thm. 4.7)
@@
LRNR approximation the entropy solution (Section 5).
@@
[…]
@@
    $\Xi_{k}$ time-layered approximation $\Xi_{k}(x,t)=\hat{I}(x)+t\,v_{k}(x)$ (5.12) for the time window $[t_{k},t_{k+1}]$
@@
    $\barwedge_{k}$ spatial discretization of the time-layered approximation $\Xi_{k}$ for the time window $[t_{k},t_{k+1}]$ satisfying
@@
    $\bar{h}$ a continuous piecewise linear approximation $\bar{h}(x,t):=\hat{u}_{0\varepsilon}\circ^{+}\doublebarwedge(x,t)$ (5.32) satisfying
@@ Appendix B Proofs of theorems and lemmas
for some coefficients $\bar{c}_{k}(\beta)$ that depend linearly on $\beta$, and $\mathcal{X}:=\{x_{k}\}_{k=1}^{K}=\cup_{i}\{x_{ij}\}_{j=1}^{N_{i}}$. Recall that $g(\alpha)=\sum_{i=1}^{r}\alpha_{i}\phi_{i}.$ Then, for corresponding continuous piecewise linear $f_{\varepsilon}(\beta)$ and $g_{\varepsilon}(\alpha)$, we have by Lemma 4.3,
@@
We reorganize the NNs as 2-layer LRNR $H_{r}$ in the form (4.16) in Example 4.2. Let $\{e_{i}\}_{i=1}^{r}$ denote the standard basis of $\mathbb{R}^{r}$. We write the weights $W^{1},W^{2}$ as
@@
(B.16) $\mathbb{U}_{r}^{1}$ $=(U_{i}^{1})_{i=1}^{r},$ $\left[U_{i}^{1}\right]_{j}:=1,\text{ for all }i,j,$
$\mathbb{U}_{r}^{2}$ $=(U_{i}^{2})_{i=1}^{r},$$\begin{cases}\left[U_{i}^{2}\right]_{j}:=\bar{c}_{\lfloor j/2\rfloor}(e_{i})&\text{ for }i=1,...\,,r-1\\ \left[U_{i}^{2}\right]_{j}:=0&\text{ for }i=r.\end{cases}$
@@
(B.18) $\mathbb{V}_{r}^{1}$ $=(V_{i}^{1})_{i=1}^{r},$$\begin{cases}\left[V_{i}^{1}\right]_{j}:=\psi_{i}(x_{\lfloor j/2\rfloor})&\text{ for }i=1,...\,,r-1,\\ \left[V_{i}^{1}\right]_{j}:=0&\text{ for }i=r,\end{cases}$
$\mathbb{V}_{r}^{2}$ $=(V_{i}^{2})_{i=1}^{r},$ $V_{i}^{1}:=0,\quad\text{ for all }i.$
@@
the LRNR $h(\gamma,\theta)$ is defined as in Example 4.2 with the coefficients (B.19) and weights and biases (LABEL:eq:W1W2,eq:B1B2). By construction, it satisfies
@@
(B.25) $E_{k}(t;x_{0})$ $:=\sup_{x\in[{\vdash}(x_{0},t),{\dashv}(x_{0},t)]}\lVert{\widehat{X}(x,t)-\Xi_{k}(x,t)}\rVert$
@@
(B.26) $\overline{u}_{\text{loc}}$ $:=\sup_{\begin{subarray}{c}z\in[{\vdash}(x_{0},t),{\vdash}(x_{0},t_{k})]\\ \qquad\cup[{\dashv}(x_{0},t_{k}),{\dashv}(x_{0},t)]\end{subarray}}u_{0}(z),\quad$
$\underline{u}_{\text{loc}}$ $:=\inf_{\begin{subarray}{c}z\in[{\vdash}(x_{0},t),{\vdash}(x_{0},t_{k})]\\ \qquad\cup[{\dashv}(x_{0},t_{k}),{\dashv}(x_{0},t)]\end{subarray}}u_{0}(z).$
Since $[\overline{u}_{\text{loc}},\underline{u}_{\text{loc}}]\subset u_{0}(\Omega_{x})$, we obtain the first estimate (5.14).
@@
then it is clear that the set of solutions forms a transported subspace, so Theorem 4.5 applies. Recall that the approximation proceeds below: Letting $\widehat{\Omega}_{x}=\Omega_{x}$, define $u_{0\varepsilon},X_{\varepsilon}\in C(\Omega_{x})\cap P_{1}(\Omega_{x})$ for which Lemma 4.3 implies
@@
then picking $X_{\varepsilon}$ so that it interpolates $X$ at the points in $\mathcal{X}\subset\Omega_{x}$ as set in Lemma 4.3, the norm $\lVert{X-X_{\varepsilon}}\rVert_{L^{\infty}(\mathcal{X})}=0$, so we have
@@
As shown in (B.20) in the proof of Theorem 4.5, there is a LRNR $h$ satisfying
@@
The coefficients $\gamma,\theta$ are induced by the mapping $\mu$ given by Theorem 4.5, see (B.19). The transported subspace coefficients depend linearly on $t$ or are constant respect to it, thus $\gamma,\theta$ are linear or constant functions of $t$.
B.6 Proof of Lemma 5.3
By construction, $\doublebarwedge(\cdot,t)$ is based on $\eta_{k}\in C(\widehat{\Omega}_{x})\cap P_{1}(\widehat{\Omega}_{x})$ given in (5.22) for index $k=1,...\,,K-1$ in (5.26) so we have for $t\in[t_{k},t_{k+1}]$
@@
upon applying (5.23) and (5.25). The RHS does not depend on the interval $[t_{k},t_{k+1}]$ that $t$ belongs to, so the result follows.
@@
We initially suppose $u_{0}\in\overline{\mathcal{U}}$ so that $u_{0}=\bar{u}_{0}$. We begin by providing an upper bound on the derivative of $X$. Recall that $X$ is continuous and piecewise smooth, therefore its weak derivative $\widehat{X}^{\prime}$ is in $L^{\infty}(\widehat{\Omega}_{x})$. Decomposing the extended domain into the shock set $\Upsilon_{t}$ and its complement
@@
Due to Lemma 4.3, we have
@@
(B.43) $\lVert{u(\cdot,t)-\bar{h}(\cdot,t))}\rVert_{L^{1}(\Omega_{x})}$
$\quad\leq\frac{1}{K}\left\lvert{u_{0}}\right\rvert_{TV}\left(4+2T\lVert{F^{\prime\prime}}\rVert_{L^{\infty}(u_{0}(\Omega_{x}))}\right)+\frac{11}{K}\left\lvert{u_{0}}\right\rvert_{TV}^{2}T\lVert{F^{\prime\prime}}\rVert_{L^{\infty}(u_{0}(\Omega_{x}))}.$
@@
$\leq\frac{11}{K}\left\lvert{u_{0}}\right\rvert_{TV}\left(1+\left\lvert{u_{0}}\right\rvert_{TV}\right)\left(1+T\lVert{F^{\prime\prime}}\rVert_{L^{\infty}(u_{0}(\Omega_{x}))}\right).$
@@
and the initial input is $Z^{0}:=x$. We remind the reader that in our setting, the spatial variable $x$ is the sole input to the NN, and the time variable $t$ is an external parameter. We use the notation $[N]:=\{1,2,...\,,N\}$ for $N\in\mathbb{N}$ and the Einstein summation convention for the subscripts. We will use higher-dimensional tensors to specify the hidden variables and the affine mappings instead of using vectors and matrices, in order to simplify the index notation.
@@
Then we have that $h(x,t)$ is equal to $\bar{h}(x,t)$, so the error estimate follows from that for $\bar{h}$ (Theorem 5.5). All the weights and biases can be written as a linear combination of at most two basis elements, and the coefficients are either constant or linear with respect to $t$.

### v4

@@ 5.2 A time-layered approximation of the rarefied characteristics
(5.7) ${\vdash},\,{\dashv}:\{(x,t)\mid x\in\Upsilon_{t},t\in\Omega_{t}\}\to\widehat{\Omega}_{x},$
@@
Since $\widehat{X}(\cdot,t)$ is continuous and it is constant in the interval ${\vdash}(x,t),{\dashv}(x,t)]$,
@@
(5.10) $\left({\vdash}(\widehat{\Omega}_{x},t)\right)=\#\left({\dashv}(\widehat{\Omega}_{x},t)\right)\leq\#\left(\{x\in\widehat{\Omega}_{x}:u_{0}(x_{-})\neq u_{0}(x_{+})\}\right),$
where the RHS is the number of jumps in the initial condition $u_{0}$, and $#(\cdot)$ denotes the cardinality of the set. For such an initial condition, the number of jumps does not increase [76].
@@
(5.11)  $\text{ if }\quad 0\leq t_{1}<t_{2}\leq T\quad\text{ then }\quad{\vdash}(x,t_{2})<{\vdash}(x,t_{1})\,\text{ and }\,{\dashv}(x,t_{1})<{\dashv}(x,t_{2}),$
@@
Due to the differential equations that yield this evolution (2.22), the endpoints $\vdash$$\dashv$} are piecewise continuously differentiable with respect to $t$, except at finitely many points in $\Omega_{t}$.
@@
(5.13) $v_{k}(x):=\left\{\begin{aligned} (F^{\prime}\circ\hat{u}_{0})(x)&&\text{ if }\lambda(x)>t_{k},\\ (F^{\prime}\circ\hat{u}_{0})({\vdash}(x,\,t_{k}))&&\text{ if }\lambda(x)\leq t_{k}.\end{aligned}\right$.
Note that $v_{k}(x)$ is constant in each connected component of the domain $\Upsilon_{t_{k}}=\{x\in\widehat{\Omega}_{x}:\lambda(x)\leq t_{k}\}$, and that $v_{k}$ can be defined using $\dashv$} in place of $\vdash$}. The function $\Xi_{k}$ satisfies the following estimates.
@@
Next, suppose we let $(t_{k})_{k=1}^{K}$ be a uniform grid over time domain $\Omega_{t}$,
@@
See Fig. 8 for an illustration of this approximation.
@@ 5.3 The spatial discretization of time-layered approximation
In the following, we will make use of the estimates (5.14) and (5.15) in Lem. 5.1, and the uniform time grid $t_{k}$ with grid size $K$ (5.16).
Because of the TV bound (5.15), there is an adaptive spatial grid $(z_{i})_{i=1}^{K}$ over $\widehat{\Omega}_{x}$ on which we have continuous piecewise linear approximations $\barwedge_{k}(\cdot,t)\in C(\widehat{\Omega}_{x}\times\Omega_{t})\cap P_{1}(\widehat{\Omega}_{x}\times\Omega_{t})$ of $\Xi_{k}(\cdot,t)$ satisfying
(5.19) $\lVert{\Xi_{k}(\cdot,t)-\barwedge_{k}(\cdot,t)}\rVert_{L^{\infty}(\widehat{\Omega}_{x})}\leq\frac{1}{K}\left(1+C_{u_{0},T,F}\right),\quad\text{ for }k=1,...\,,K-1.$
@@
(5.20) $\barwedge_{k}(z_{i},t_{k})=\left\{\begin{aligned} \widehat{X}_{0}(z_{i},t_{k})&\quad\text{ for }z_{i}\in\widehat{\Omega}_{x}\setminus\Upsilon_{t_{k}},\\ \widehat{X}(z_{i},t_{k})&\quad\text{ for }z_{i}\in\Upsilon_{t_{k}}.\end{aligned}\right$.
@@
(5.21) $\eta_{k}(z_{i})=\frac{1}{t_{k+1}}(\barwedge_{k+1}(z_{i},t_{k})-\hat{I}(z_{i}))-\frac{1}{t_{k}}(\barwedge_{k}(z_{i},t_{k})-\hat{I}(z_{i}))\quad i=0,1,...\,,K-1.$
Then $\eta_{k}$ is nonzero only at the grid points $(z_{i})_{i=1}^{K}$ that lie in $\Upsilon_{t_{k+1}}\cup\Upsilon_{t_{k}}$. See Fig. 9 for an illustration.
@@
Next, let us set the function $\barwedge:\widehat{\Omega}_{x}\times\Omega_{t}\to\mathbb{R}$ which is continuous piecewise linear in space but only piecewise linear in time. Let $\barwedge_{K}(x,t):=\hat{I}(x)+\frac{t}{T}(\barwedge_{K}(x,t_{K})-\hat{I}(x))$ and
@@
$\leq\lVert{\widehat{X}(\cdot,t)-\Xi_{k}(\cdot,t)}\rVert_{L^{\infty}(\widehat{\Omega}_{x})}$
$\quad~~~+\lVert{\Xi_{k}(\cdot,t)-\barwedge_{k}(\cdot,t_{k})}\rVert_{L^{\infty}(\widehat{\Omega}_{x})}+\lVert{\barwedge_{k}(\cdot,t_{k})-\barwedge(\cdot,t)}\rVert_{L^{\infty}(\widehat{\Omega}_{x})}$
@@
Figure 10: A diagram depicting the approximation by $\doublebarwedge(x,t)$. The term $F^{\prime}\circ\hat{u}_{0}$ in the rarefied characteristics with multivalued inverse $\widehat{X}_{0}$ along with the interpolant $\eta_{k}$ (5.21) are shown in the upper left plot. The interpolant $\eta_{k}$ along with the left and right shock endpoints $\vdash$$\dashv$} at times $t_{k}$ and $t_{k+1}$ are shown in the upper center plot. The three lower plots show a movie illustrating the evolution of $\doublebarwedge(x,t)$ starting at time $t=t_{k}$. The thin interpolant $\eta_{k}$ representing a tiny perturbation of the characteristic curves, is moving out of the spatial domain to the right, traveling quickly outside the domain.
@@ Appendix A Glossary of definitions and relations
Formulations of the entropy solution (Sec. 2).
@@
    $\widehat{X}_{0}$ rarefied characteristics with multivalued inverse
[…]
@@
    $\rho=\mathbf{1}_{\mathbb{R}_{+}}$ threshold function
@@
Low rank neural representation (LRNR) (Sec. 3).
@@
Transported subspaces, inverse-bias trick, LRNR approximation of classical solutions (Sec. 4).
@@
    LRNR ($L=2,r=3$) approximation $h(t)\in H_{r}$ to the classical solution (Thm. 4.8)
@@
LRNR approximation the entropy solution (Sec. 5).
@@
    $\vdash$$\dashv$} left and right shock endpoints (5.8)
@@
    $\Xi_{k}$ time-layered approximation $\Xi_{k}(x,t)=\hat{I}(x)+t\,v_{k}(x)$ (5.12) for the time window $t_{k},t_{k+1}]$
@@
    $\barwedge_{k}$ spatial discretization of the time-layered approximation $\Xi_{k}$ for the time window $t_{k},t_{k+1}]$ satisfying
@@
    $\bar{h}$ continuous piecewise linear approximation $\bar{h}(x,t):=\hat{u}_{0\varepsilon}\circ^{+}\doublebarwedge(x,t)$ (5.31) satisfying
@@ Appendix B Proofs of theorems and lemmas
for some coefficients $\bar{c}_{k}(\beta)$ that depend linearly on $\beta$, and $\mathcal{X}:=\{x_{k}\}_{k=1}^{K}=\cup_{i}\{x_{ij}\}_{j=1}^{N_{i}}$. Recall that $g(\alpha)=\sum_{i=1}^{r}\alpha_{i}\phi_{i}.$ Then, for corresponding continuous piecewise linear $f_{\varepsilon}(\beta)$ and $g_{\varepsilon}(\alpha)$, we have by Lem. 4.3,
@@
We reorganize the NNs as 2-layer LRNR $H_{r}$ in the form (4.16) in Exmpl. 4.5. Let $\{e_{i}\}_{i=1}^{r}$ denote the standard basis of $\mathbb{R}^{r}$. We write the weights $W^{1},W^{2}$ as
@@
(B.16) $\mathbb{U}_{r}^{1}$ $=(U_{i}^{1})_{i=1}^{r},$ $\left[U_{i}^{1}\right]_{j}=1,\text{ for all }i,j,$
[…]
@@
[…]
$\mathbb{V}_{r}^{2}$ $=(V_{i}^{2})_{i=1}^{r},$ $V_{i}^{1}=0,\quad\text{ for all }i.$
@@
the LRNR $h(\gamma,\theta)$ is defined as in Exmpl. 4.5 with the coefficients (B.19) and weights and biases (B.15-B.17). By construction, it satisfies
@@
(B.25) $E_{k}(t;x_{0})$ $=\sup_{x\in[{\vdash}(x_{0},t),{\dashv}(x_{0},t)]}\lVert{\widehat{X}(x,t)-\Xi_{k}(x,t)}\rVert$
@@
(B.26) $\overline{u}_{\text{loc}}$ $=\sup_{\begin{subarray}{c}z\in[{\vdash}(x_{0},t),{\vdash}(x_{0},t_{k})]\\ \hskip 16.38895pt\cup[{\dashv}(x_{0},t_{k}),{\dashv}(x_{0},t)]\end{subarray}}u_{0}(z),\quad$
$\underline{u}_{\text{loc}}$ $=\inf_{\begin{subarray}{c}z\in[{\vdash}(x_{0},t),{\vdash}(x_{0},t_{k})]\\ \hskip 16.38895pt\cup[{\dashv}(x_{0},t_{k}),{\dashv}(x_{0},t)]\end{subarray}}u_{0}(z).$
Since $\overline{u}_{\text{loc}},\underline{u}_{\text{loc}}]\subset u_{0}(\Omega_{x})$, we obtain the first estimate (5.14).
@@
B.4.1 Remark
Note that the upper bound (5.14) with the TV norm $\left\lvert{u_{0}}\right\rvert_{TV(\Omega_{x})}$ obtained from the last term in (B.25) is not sharp, and can be replaced by $2\lVert{u_{0}}\rVert_{L^{\infty}(\Omega_{x})}$ because $\left\lvert\overline{u}_{\text{loc}}-\underline{u}_{\text{loc}}\right\rvert\leq 2\lVert{u_{0}}\rVert_{L^{\infty}([{\vdash}(x_{0},t),{\dashv}(x_{0},t)])},$ but we will keep the TV norm here, as doing so this will not affect the final results of this work.
@@ B.4.1 Remark
then it is clear that the set of solutions forms a transported subspace, so Thm. 4.6 applies. Recall that the approximation proceeds below: Letting $\widehat{\Omega}_{x}=\Omega_{x}$, define $u_{0\varepsilon},X_{\varepsilon}\in C(\Omega_{x})\cap P_{1}(\Omega_{x})$ for which Lem. 4.3 implies
@@
then picking $X_{\varepsilon}$ so that it interpolates $X$ at the points in $\mathcal{X}\subset\Omega_{x}$ as set in Lem. 4.3, the norm $\lVert{X-X_{\varepsilon}}\rVert_{L^{\infty}(\mathcal{X})}=0$, so we have
@@
As shown in (B.20) in the proof of Thm. 4.6, there is a LRNR $h$ satisfying
@@
The coefficients $\gamma,\theta$ are induced by the mapping $\mu$ given by Thm. 4.6, see (B.19). The transported subspace coefficients depend linearly on $t$ or are constant respect to it, thus $\gamma,\theta$ are linear or constant functions of $t$.
B.6 Proof of Theorem 5.3
By construction, $\doublebarwedge(\cdot,t)$ is based on $\eta_{k}\in C(\widehat{\Omega}_{x})\cap P_{1}(\widehat{\Omega}_{x})$ given in (5.21) for index k=1,,K1k=1,...\,,K-1 in (5.25) so we have for $t\in[t_{k},t_{k+1}]$
@@
upon applying (5.22) and (5.24). The RHS does not depend on the interval $t_{k},t_{k+1}]$ that $t$ belongs to, so the result follows.
@@
We initially suppose $u_{0}\in\overline{\mathcal{U}}$ so that $u_{0}=\bar{u}_{0}$. We begin by providing an upper bound on the derivative of $\widehat{X}$. Recall that $\widehat{X}$ is continuous and piecewise smooth, therefore its weak derivative $\widehat{X}^{\prime}$ is in $L^{\infty}(\widehat{\Omega}_{x})$. Decomposing the extended domain into the shock set $\Upsilon_{t}$ and its complement
@@
Due to Lem. 4.3, and choosing $\epsilon=1/K$ so that $\lVert{\rho-\rho_{\varepsilon}}\rVert_{L^{1}(\mathbb{R})}\lesssim 1/K$ we have
@@
(B.43) $\lVert{u(\cdot,t)-\bar{h}(\cdot,t))}\rVert_{L^{1}(\Omega_{x})}\lesssim\frac{1}{K}\left\lvert{u_{0}}\right\rvert_{TV}+\frac{1}{K}\left\lvert{u_{0}}\right\rvert_{TV}^{2}T\lVert{F^{\prime\prime}}\rVert_{L^{\infty}(u_{0}(\Omega_{x}))}.$
@@
$\lesssim\frac{1}{K}\left\lvert{u_{0}}\right\rvert_{TV}\left(1+\left\lvert{u_{0}}\right\rvert_{TV}\right)\left(1+T\lVert{F^{\prime\prime}}\rVert_{L^{\infty}(u_{0}(\Omega_{x}))}\right).$
@@
and the initial input is $Z^{0}:=x$. We remind the reader that in our setting, the spatial variable $x$ is the sole input to the NN, and the time variable $t$ is an external parameter. We use the notation $N]:=\{1,2,...\,,N\}$ for $N\in\mathbb{N}$ and the Einstein summation convention for the subscripts. We will use higher-dimensional tensors to specify the hidden variables and the affine mappings instead of using vectors and matrices, in order to simplify the index notation.
@@
Then we have that $h(x,t)$ is equal to $\bar{h}(x,t)$, so the error estimate follows from that for $\bar{h}$ (Thm. 5.5). All the weights and biases can be written as a linear combination of at most two basis elements, and the coefficients are either constant or linear with respect to $t$.